beamline experiments
  Beamline Logbook  Logged in as "Eric Dufresne" ELOG logo
Message ID: 75     Entry time: Mon Jun 21 17:41:01 2004     In reply to: 73
Author: Eric Dufresne 
Type: Routine 
Category: BPM 
Subject: time series from 06/18 to 06/21 in top-up hybrid mode 
  
Here is data from the weekend started at 17h02 on Friday 6/18 and lasting
about 64 hrs until 6/21 at 8am when the beam dumps. There were no top up failures
during this time.
 
Attachment 1 shows the intensities. Experiments are ungoing so access  to
7ID-C sends the BPM values to large negative values. Linda's group scanned
the monochromator energy several times(note that it moves the beam in the
vertical), detuned the mono 2nd crystal several times. 
  
Attachment 2 shows the 7ID-C X-ray BPM data. Note that the vertical beam moves by
about 75 um peak to peak during energy scans, so one expect the transmitted
intensity to vary some.
 
Attachment 3 shows the APS XBPM data projected to the L5-20 plane. The APS beam
position was very stable.
 
Attachment 4 shows the cryocooler information.
 
Attachment 5 shows the piezo voltage, showing when several tweaks happened during
several energy scans of the mono.
 
Eric D
   
Attachment 1: 061804_17h02m19s_int.png  8 kB
Attachment 2: 061804_17h02m19s_bpm.png  8 kB
Attachment 3: 061804_17h02m19s_apsbpm.png  6 kB
Attachment 4: 061804_17h02m19s_mono.png  7 kB
Attachment 5: 061804_17h02m19s_piezo.png  5 kB
ELOG V2.5.2